slipknot3066:

Solve. {d+e=2d−e=4 Use the substitution method. There are an infinite number of solutions. The solution is (5, 1). There is no solution. The solution is (3, −1).

sillybilly123:

\(d+e=2 \qquad \square\) \(d - e = 4 \qquad \triangle\) from \(\triangle\): \(d = e + 4\) Pop that into \(\square\): \(e + 4 +e=2\) And solve it for e, then solve for d